Consider the following data assuming that these are the only products produced and consumed in the economy: Product 2017 Quantity 2017 Price ($) 2018 Quantity 2018 Price ($) Beer 14 18 16 13 Chocolate 15 11 17 12 Using 2018 as the base year answer the following questions: a. What was the nominal GDP in 2017 and 2018? b. What was the real GDP in 2017 and 2018? c. What was the GDP deflator in 2017 and 2018? d. What was the growth rate of nominal GDP between 2017 and 2018? What was the growth rate of real GDP between 2017 and 2018? What was the growth rate of the GDP deflator between 2017 and 2018? e. Calculate the value of the basket in 2017 and 2018. f. Calculate the CPI for 2017 and 2018. g. Calculate the annual inflation rate over the 2017-2018 period. h. (This part is OPTIONAL in case you want more practice, this type of question is usually on the final exam :-) Change your base year from 2018 to 2017. Re-do your answers to parts a, b, and c.
Answer #1
Product Q2017 P2017 Q2018 P2018
Beer 14 18 16 13
Chocolate 15 11 17 12

a.

Nominal GDP is sum of quantity multiplied to price of all goods and services produced in the year.

NGDP2017 = 14*18+15*11 = 417

NGDP2018 = 16*13+17*12 = 412

b.

Real GDP is sum of quantity multiplied to price of all goods and services as per base year, produced in the year. The base year is 2018.

RGDP2017 = 14*13+15*12 = 362

RGDP2018 = 16*13+17*12 = 412

c.

GDP deflator is NGDP divided by RGDP multiplied by 100

GDP deflator 2017 = 417/362*100 = 115.19

GDP deflator 2018 = 412/412*100 = 100

d.

The nominal GDP growth rate between 2017-2018 is = (412-417)/417*100 = -1.2%